Farmer received rupees 783000 after deducting his marketing cost of rupees 6 per quintal by selling 750 quintals of paddy. find out the total marketing cost incurred by the farmer. The wholesaler bought the farmers produce at rupees 1050 per quintal incurred Rs.48.937/ quintal sold at Rs.1140 per quintal. Find out his margin?similarly the retailer body produced from the wholesaler and incurred Rs.15.3 63 per quintal and sold to consumer at Rs.1200 per quintal now find out the margin of retailer and the producers share in consumers rupee.
1
Expert's answer
2020-12-17T07:39:28-0500

The total marketing cost incurred by the farmer is: TC = 750*6 = 4500.

The wholesaler margin is: "(1140 - 1050)\/1140 = 0.079" or 7.9%.

The retailer body margin is: "(1200 - 1140)\/1200 = 0.05" or 5%.